Calculate ∆H for the reaction CH4 (g) + NH3 (g) → HCN (g) + 3 H2 (g), from the reactions.

N2 (g) + 3 H2 (g) → 2 NH3 (g) ∆H = -91.8 kJ

C (s, graphite) + 2 H2 (g) → CH4 (g) ∆H = -74.9 kJ/mole

H2 (g) + 2 C (s, graphite) + N2 (g) → 2 HCN (g) ∆H = +270.3 kJ


1
Expert's answer
2022-06-01T03:14:04-0400

eq - equation

(1st eq): N2(g) + 3H2(g) → 2NH3(g), ∆H1 = −91.8 kJ

(2nd eq): C(s, graphite) + 2H2(g) → CH4(g), ∆H2 = −74.9 kJ/mol

(3d eq): H2(g) + 2C(s, graphite) + N2(g) → 2HCN(g), ∆H3 = +270.3 kJ


CH4(g) + NH3(g) → HCN(g) + 3H2(g), ∆Hx = ???


Solution:

According to Hess's law, the heat of reaction depends upon Initial and final conditions of reactants and does not depend of the intermediate path of the reaction.


1) Modify the three given equations to get the target equation:

(1st eq): flip it over and divide it by 2. We need 1 mol of NH3(g) on the reactant side.

(2nd eq): flip it over. We need 1 mol of CH4(g) on the reactant side.

(3d eq): divide it by 2. We need 1 mol of HCN(g) on the product side.


2) Rewrite the three given equations with the changes made (including changes in their enthalpy):

(1*st eq): NH3(g) → 0.5N2(g) + 1.5H2(g), ∆H1* = (−91.8 kJ) × (−1/2) = +45.9 kJ

(2*nd eq): CH4(g) → C(s, graphite) + 2H2(g), ∆H2* = (−74.9 kJ) × (−1) = +74.9 kJ

(3*d eq): 0.5H2(g) + C(s, graphite) + 0.5N2(g) → HCN(g), ∆H3* = (+270.3 kJ) × (1/2) = +135.15 kJ


3) Cancel out the common species on both sides:

0.5N2(g) ⇒ (3*d equation) & (1*st equation)

C(s, graphite) ⇒ (3*d equation) & (2*nd equation)

0.5H2(g) ⇒ (3*d equation) & (sum of 1*st and 2*nd equations)


Thus, adding modified equations and canceling out the common species on both sides, we get:

CH4(g) + NH3(g) → HCN(g) + 3H2(g)


3) Add the ΔH values of (1*st), (2*nd) and (3*d eq) equations to get your answer:

∆Hx = ∆H1* + ∆H2* + ∆H3*

∆Hx = (+45.9 kJ) + (+74.9 kJ) + (+135.15 kJ) = +255.95 kJ

∆Hx = +255.95 kJ


Answer: ∆H for the reaction is +255.95 kJ